OpenStudy (anonymous):

Write the composite function in the form f(g(x)). [Identify the inner function u = g(x) and the outer function y = f(u). y=e^7 rootx would it be (g(x),f(u))= (e^x, 7rootx)?

OpenStudy (freckles):

\[y=e^{\sqrt[7]{x}}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait n, ill draw it

OpenStudy (freckles):

it looks like you have identified things a bit backwards

OpenStudy (anonymous):



OpenStudy (anonymous):

the 7 and the square root are being multiplied

OpenStudy (freckles):

\[y=e^{7 \sqrt{x}}\] you mean this?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

say we have \[f(x)=e^x\] what would you get if I asked you to find \[f(7 \sqrt{x})=?\]

OpenStudy (freckles):

\[f( \text{ twinkie }-\text{ star})=e^{\text{ twinkie}-\text{ star}}\]

OpenStudy (freckles):

just giving you example of how to do what I asked if you were having trouble

OpenStudy (anonymous):

would it be (7-e^x)=e^(7-e^x)?

OpenStudy (freckles):